OpenStudy (harleyquinn):

Can someone help me find the X and Y intercepts? 3x+6y=-24

OpenStudy (anonymous):

y intercept is where x=0, so y = -24/6 = -4 x intercept is where y=0, so x = -24/3 = -8

OpenStudy (harleyquinn):

Okay so, you let x=0 to find Y? And y=0 to find X?

OpenStudy (anonymous):

use for multiple choice http://www.mathway.com/problem.aspx?p=precalculus totally not cheating my teacher form flvs gave it to me

OpenStudy (anonymous):

yup, think about it graphically.

OpenStudy (harleyquinn):

Thanks rain_ I go to flvs too. But I need to know how to do it. I have a test tomorrow. And thanks alexnhchong.

OpenStudy (anonymous):

kk